# Sifan Wang Graduate Group in Applied Mathematics and Computational Science

### 01 Feb 2024 v1

@@ 5 Results
Ginzburg-Laudau $\mathbf{1.49\times 10^{-2}}$ $3.20\times 10^{-2}$
@@ 5.1 Allen-Cahn equation
Original formulation of Raissi et al. [54] $014.98e-$ 01
Adaptive time sampling [49] $022.33e-$ 02
Self-attention [31] $022.10e-$ 02
Time marching [74] $021.68e-$ 02
Causal training [30] $041.39e-$ 04
Dirac delta function causal training [60] $056.29e-$ 05
JAX-PI [62] $055.37e-$ 05
RBA-PINNs [61] $054.55e-$ 05
PirateNet (Ours) $\mathbf{2.24e-05}$
@@ 5.2 Korteweg–De Vries equation
Dirac delta function causal training [60] $032.45e-$ 03
JAX-PI [62] $031.96e-$ 03
PirateNet (Ours) $\mathbf{4.27e-04}$
@@ 5.3 Grey-Scott equation
5.4 Ginzburg-Laudau equation
@@
Figure 11: Ginzburg-Laudau equation: Comparisons between the solutions predicted by a trained PirateNet and the reference solutions. The detailed hyper-parameter settings are presented in Table 7.
Figure 12: Ginzburg-Laudau equation: Convergence of the initial condition losses, the PDE residual losses and the relative $L^{2}$ errors during the training of a PirateNet and a Modified MLP backbone with different activation functions.
Figure 13: Ginzburg-Laudau equation: Left: Relative $L^{2}$ errors of training PirateNet of different depth, with the nonlinearity parameter $\alpha$ of each residual block initialized to 0 or 1. Right: Relative $L^{2}$ errors of training PirateNet with and without gating operations, respectively. Each ablation study is performed under the same hyper-parameter setting. All models are trained within the first time window, with the results averaged over five random seeds.
@@ 5.5 Lid-driven Cavity flow
where $x\in[0,1],y=1,C_{0}=50$. For the other three walls, we enforce a no-slip boundary condition. Our goal is to obtain the velocity and pressure filed corresponding a Reynolds number of $3,200$.
@@
Figure 14 plots the predicted velocity field for a Reynolds number of $Re=3,200$. The predictions align closely with the reference results from Ghia et al. [78]. This is evidenced by a relatively low $L^{2}$ error of $4.21\times 10^{-2}$, a significant improvement over the $1.58\times 10^{-1}$ error reported by JAXPI [62]. Our ablation study on the model’s depth is visualized on the right panel of Figure 15, which reveals consistent performance enhancements as the depth increases within a moderately large range. However, the model seems to be saturated at depths greater than 20, yet the training remains stable and produces results similar to the one at a depth of 18. This observation highlights PirateNet’s impressive scalability and stability in handling complex, highly nonlinear problems and motivates future advancements in the model architecture. Interestingly, we also examine the evolution of the nonlinearities $\alpha$ in each residual block of the network. A noticeable increase in $\alpha$ is observed when the model transitions to training at higher Reynolds numbers. This trend suggests that the model requires a higher degree of nonlinearity in order to minimize the associated PDE residuals at higher Reynolds numbers.
@@
Figure 15: Lid-driven cavity (Re=$3,200$): Left: Convergence of the relative $L^{2}$ errors during the training of PirateNets of different depth for the final training stage at $Re=3,200$. All experiments are performed under the same hyper-parameter setting. Right: Evolution of nonlinearities in each PirateNet residual block of training a PirateNet with 6 residual blocks (18 hidden layers).

### 05 Feb 2024 v2

@@ 5 Results
Ginzburg-Landau $\mathbf{1.49\times 10^{-2}}$ $3.20\times 10^{-2}$
@@ 5.1 Allen-Cahn equation
Original formulation of Raissi et al. [54] $4.98\times 10^{-1}$
Adaptive time sampling [49] $2.33\times 10^{-2}$
Self-attention [31] $2.10\times 10^{-2}$
Time marching [74] $1.68\times 10^{-2}$
Causal training [30] $1.39\times 10^{-4}$
Dirac delta function causal training [60] $6.29\times 10^{-5}$
JAX-PI [62] $5.37\times 10^{-5}$
RBA-PINNs [61] $4.55\times 10^{-5}$
PirateNet (Ours) $\mathbf{2.24\times 10^{-5}}$
@@ 5.2 Korteweg–De Vries equation
Dirac delta function causal training [60] $2.45\times 10^{-3}$
JAX-PI [62] $1.96\times 10^{-3}$
PirateNet (Ours) $\mathbf{4.27\times 10^{-4}}$
@@ 5.3 Grey-Scott equation
5.4 Ginzburg-Landau equation
@@
Figure 11: Ginzburg-Landau equation: Comparisons between the solutions predicted by a trained PirateNet and the reference solutions. The detailed hyper-parameter settings are presented in Table 7.
Figure 12: Ginzburg-Landau equation: Convergence of the initial condition losses, the PDE residual losses and the relative $L^{2}$ errors during the training of a PirateNet and a Modified MLP backbone with different activation functions.
Figure 13: Ginzburg-Landau equation: Left: Relative $L^{2}$ errors of training PirateNet of different depth, with the nonlinearity parameter $\alpha$ of each residual block initialized to 0 or 1. Right: Relative $L^{2}$ errors of training PirateNet with and without gating operations, respectively. Each ablation study is performed under the same hyper-parameter setting. All models are trained within the first time window, with the results averaged over five random seeds.
@@ 5.5 Lid-driven Cavity flow
where $x\in[0,1],y=1,C_{0}=50$. For the other three walls, we enforce a no-slip boundary condition. Our goal is to obtain the velocity and pressure filed corresponding a Reynolds number of $3200$.
@@
Figure 14 plots the predicted velocity field for a Reynolds number of $Re=3200$. The predictions align closely with the reference results from Ghia et al. [78]. This is evidenced by a relatively low $L^{2}$ error of $4.21\times 10^{-2}$, a significant improvement over the $1.58\times 10^{-1}$ error reported by JAXPI [62]. Our ablation study on the model’s depth is visualized on the right panel of Figure 15, which reveals consistent performance enhancements as the depth increases within a moderately large range. However, the model seems to be saturated at depths greater than 20, yet the training remains stable and produces results similar to the one at a depth of 18. This observation highlights PirateNet’s impressive scalability and stability in handling complex, highly nonlinear problems and motivates future advancements in the model architecture. Interestingly, we also examine the evolution of the nonlinearities $\alpha$ in each residual block of the network. A noticeable increase in $\alpha$ is observed when the model transitions to training at higher Reynolds numbers. This trend suggests that the model requires a higher degree of nonlinearity in order to minimize the associated PDE residuals at higher Reynolds numbers.
@@
Figure 15: Lid-driven cavity (Re=$3200$): Left: Convergence of the relative $L^{2}$ errors during the training of PirateNets of different depth for the final training stage at $Re=3200$. All experiments are performed under the same hyper-parameter setting. Right: Evolution of nonlinearities in each PirateNet residual block of training a PirateNet with 6 residual blocks (18 hidden layers).